OpenStudy (anonymous):

check your common sense if, 1=5 2=25 3=325 4=4325 then 5=?

OpenStudy (anonymous):

5= 54325 ?

OpenStudy (anonymous):

I got that answer from a website heres the explanation. 1 = "5" =5 '2' = '2',"5" =25 -3- = -3-'2',"5" =325 +4+ = +4+,-3-,'2',"5" =4325 .5. = .5.,+4+,-3-,'2',"5" =54325

OpenStudy (anonymous):

O.o looks complicated

OpenStudy (anonymous):

5325 :P